# TimeDistill: Efficient Long-Term Time Series Forecasting with MLP via Cross-Architecture Distillation

### v1

@@
Abstract
Transformer-based and CNN-based methods demonstrate strong performance in long-term time series forecasting. However, their high computational and storage requirements can hinder large-scale deployment. To address this limitation, we propose integrating lightweight MLP with advanced architectures using knowledge distillation (KD). Our preliminary study reveals different models can capture complementary patterns, particularly multi-scale and multi-period patterns in the temporal and frequency domains. Based on this observation, we introduce TimeDistill, a cross-architecture KD framework that transfers these patterns from teacher models (e.g., Transformers, CNNs) to MLP. Additionally, we provide a theoretical analysis, demonstrating that our KD approach can be interpreted as a specialized form of mixup data augmentation. TimeDistill improves MLP performance by up to 18.6%, surpassing teacher models on eight datasets. It also achieves up to 7$\times$ faster inference and requires 130$\times$ fewer parameters. Furthermore, we conduct extensive evaluations to highlight the versatility and effectiveness of TimeDistill.
Machine Learning, ICML
1 Introduction
Forecasting is a notably critical problem in the time series analysis community, which aims to predict future time series based on historical time series records (Wang et al., 2024b). It has broad practical applications such as climate modeling (Wu et al., 2023), traffic flow management (Yin et al., 2021), healthcare monitoring (Kaushik et al., 2020) and finance analytics (Granger & Newbold, 2014).
@@ 1 Introduction
Figure 1: Performance comparison.
rise of Transformers in various domains (Devlin et al., 2018; Khan et al., 2022) has led to their wide adoption in time series forecasting (Wen et al., 2022; Wu et al., 2021; Zhou et al., 2022, 2021; Nie et al., 2023; Liu et al., 2024), leveraging the strong capabilities of capturing pairwise dependencies and extracting multi-level representations within sequential data. Similarly, CNN architectures have also proven effective by developing convolution blocks for time series (Luo & Wang, 2024; Wang et al., 2023). However, despite the strong performance of Transformer-based and CNN-based models, they face significant challenges in large-scale industrial applications due to their relatively high computational demands, especially in latency-sensitive scenarios like financial prediction and healthcare monitoring (Granger & Newbold, 2014; Kaushik et al., 2020). In contrast, simpler linear or MLP models offer greater efficiency, although with lower performance (Zeng et al., 2023; Lin et al., 2024). These contrasting observations raises an intriguing question:
@@
A promising approach to addressing this question is knowledge distillation (KD(Hinton, 2015), a technique that transfers knowledge from a larger and more complex model (teacher) to a smaller and simpler one (student) while maintaining comparable performance. In this work, we pioneer cross-architecture KD in time series forecasting, with MLP as the student and other advanced architectures (e.g., Transformers and CNNs) as the teacher. However, designing such a framework is non-trivial, as it remains unclear what “knowledge” should be distilled into MLP.
@@
To investigate this potential, we conduct a comparative analysis of prediction patterns between MLP and other time series models. Our findings reveal that MLP still excels on some data subsets despite its overall lower performance (Sec. 3.1), which highlights the value of harnessing the complementary capabilities across different architectures. To further explore the specific properties to distill, we focus on two key time series patterns: multi-scale pattern in temporal domain and multi-period pattern in frequency domain, given that they are vital in capturing the complex structures typical of many time series. (1) Multi-Scale Pattern: Real-world time series often show variations at multiple temporal scales. For example, hourly recorded traffic flow data capture changes within each day, while daily sampled data lose fine-grained details but reveal patterns related to holidays (Wang et al., 2024a). We observe that models that perform well on the finest scale also perform accurately on coarser scales, while MLP fails on most scales (Sec. 3.2). (2) Multi-Period Pattern: Time series often exhibit multiple periodicities. For instance, weather measurements may have both daily and yearly cycles, while electricity consumption data may show weekly and quarterly cycles (Wu et al., 2022). We find that models that perform well can capture periodicities similar to those in the ground truth, but MLP fails to capture these periodicities (Sec. 3.2). Therefore, enhancing MLP requires distilling and integrating these multi-scale and multi-period patterns from teacher models.
@@
Based on our observations, we propose a cross-architecture KD framework named TimeDistill to bridge the performance and efficiency gap between complex teacher models and a simple MLP. Instead of solely matching predictions in conventional KD, TimeDistill focuses on aligning multi-scales and multi-period patterns between MLP and the teacher: we downsample the time series for temporal multi-scale alignment and apply Fast Fourier Transform (FFT) to align period distributions in the frequency domain. The KD process can be conducted offline, shifting heavy computations from the latency-critical inference phase, where millisecond matter, to the less time-sensitive training phase, where longer processing time is acceptable. We validate the effectiveness of TimeDistill both theoretically and empirically and summarize our contributions as follows:
@@
Figure 2: Model efficiency comparison averaged across all prediction lengths (96, 192, 336, 720) for the ECL dataset. Full results on more datasets are listed in Appendix I.
@@
    We provide theoretical insights into the benefits of TimeDistill, illustrating that the proposed distillation process can be viewed as a form of data augmentation through a special mixup strategy.
@@
    We show that TimeDistill is both effective and efficient, consistently outperforming standalone MLP by up to 18.6% and surpassing teacher models in nearly all cases (see Figure 1). Additionally, it achieves up to 7x faster inference and requires up to 130× fewer parameters compared to teacher models (see Figure 2).
@@
2 Related Work
We present a brief summary of related work, with a detailed version provided in Appendix A. A Debate in Long-Term Time Series Forecasting: The trade-off between performance and efficiency has sparked debate between advanced architecture and MLP in long-term time series forecasting (Zeng et al., 2023; Zhang et al., 2022; Lin et al., 2024). Transformers (Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022; Nie et al., 2023; Liu et al., 2024) or CNNs (Luo & Wang, 2024; Wu et al., 2022; Wang et al., 2023) excel but face competition from a simple Linear or MLP model, which achieves comparable performance with greater efficiency (Zeng et al., 2023; Oreshkin et al., 2020; Challu et al., 2023; Zhang et al., 2022; Chen et al., 2023). To combine these strengths, we propose cross-architecture knowledge distillation (KD) to achieve both high performance and efficiency. KD in Time Series: CAKD (Xu et al., 2022) uses adversarial and contrastive learning for feature distillation without specific design for time series, while LightTS (Campos et al., 2023) designs a KD framework for ensemble classifiers, limiting its generality. Unlike these, our framework targets time series-specific patterns, such as multi-scale and multi-period, pioneering cross-architecture KD.
3 Preliminary Studies
@@
Figure 3: Win ratio (%) of MLP v.s. teacher models across datasets under input-720-predict-96 setting. The win ratio is generally large (average: 49.92%, median: 49.96%), indicating MLP and teacher models excel on different samples with minimal overlap.
@@
Figure 4: Visualization of model predictions on different downsampled scales of ECL dataset. MLP consistently shows poor performance at multiple scales, while other models perform well, highlighting the importance of capturing multi-scale patterns.
@@
Figure 5: Prediction spectrograms of various models on ECL dataset against the ground truth. MLP fails to match the amplitudes of several main frequencies in the ground truth, with red numbers indicating amplitude differences for the most significant frequency.
@@
3.1 What Motivates Distillation?
@@
While MLP excel in efficiency, they often lag in performance compared to Transformer and CNN models or achieve only similar results (Figure 1). However, even if MLP underperforms its teachers overall, it may excel on specific samples. To investigate this, we compare prediction errors of MLP’s prediction $\mathbf{\hat{Y}}_{s}$ and teacher model’s prediction $\mathbf{\hat{Y}}_{t}$ for $N$ samples, defined as $e_{s}=\{\text{MSE}(\mathbf{\hat{Y}}_{s}^{1},\mathbf{Y}^{1}),\cdots,\text{MSE}(\mathbf{\hat{Y}}_{s}^{N},\mathbf{Y}^{N})\}$ and $e_{t}=\{\text{MSE}(\mathbf{\hat{Y}}_{t}^{1},\mathbf{Y}^{1}),\cdots,\text{MSE}(\mathbf{\hat{Y}}_{t}^{N},\mathbf{Y}^{N})\}.$ We calculate the win ratio, indicating where MLP outperforms teachers:
Win $\text{Win Ratio}={\sum\mathbbm{1}(e_{s}<e_{t})}/{N},$ (1)
@@
where $\mathbbm{1}(\cdot)$ equals 1 if MLP outperforms the teacher, otherwise 0. As shown in Figure 5, the win ratio is high (average: 49.92%, median: 49.96%), despite MLP underperforming teacher models overall. For example, on the Traffic dataset, MLP lags behind ModernTCN but wins on 81.19% of the samples, indicating differing strengths on distinct subsets. This highlights the potential of distilling the complementary knowledge from the teacher model into MLP.
@@
An intuitive strategy of KD for MLP is to align predictions with teachers, but this faces limitations. First, it risks overfitting noise in the teacher’s predictions, leading to less stable knowledge (Chen et al., 2017; Takamoto et al., 2020; Gajbhiye et al., 2021). Second, MLP may struggle to replicate the complex patterns such as seasonality, trends, and periodicity in teacher predictions directly due to their limited capacity. Third, this approach overlooks valuable knowledge from intermediate features of teacher model. Therefore, the specific knowledge to distill into MLP requires further exploration.
3.2 What Should We Distill?
@@
To investigate which complementary time series patterns to distill into MLP, we analyze prediction patterns in both temporal and frequency domains, considering real-world variations across temporal scales (Wang et al., 2024a) and periodicities (Wu et al., 2022). Thus, we conduct further analysis of these two patterns by presenting illustrative cases. The implementation is detailed in Appendix D. (1) Multi-Scale Pattern: By downsampling predictions $\mathbf{\hat{Y}}$ using convolutional operations, we obtain their multi-scale representations. Figure 5 shows the analysis of multi-scale temporal patterns by downsampling the time series (Scale 0) to coarser scales (Scales 1–3). Models excelling at Scale 0 generally perform well across all scales. The teacher models align closely with the ground truth at Scale 3, capturing underlying trends effectively. In contrast, MLP significantly deviates, showing its limitations in handling multi-scale patterns. (2) Multi-Period Pattern: Periodicities in time series are visible in the frequency domain by transforming time series into spectrograms, with the x-axis representing frequency and periodicity calculated as the time series length $S$ divided by frequency. Figure 5 shows spectrograms of model predictions and the ground truth. Models with lower MSE capture periodicities more accurately, closely matching the ground truth at dominant frequencies. In contrast, MLP predictions show larger discrepancies, highlighting its inability to effectively capture periodic patterns.
@@
From our observations, we conjecture that MLP underperforms on certain samples due to their inability to capture essential multi-scale and multi-period patterns, which are essential for effective time series forecasting. To improve MLP, it is crucial to incorporate these complementary patterns from teacher models during distillation.
@@
Figure 6: Overall framework of TimeDistill, which distills knowledge from a teacher model to a student MLP using (a) Multi-Scale Distillation and (b) Multi-Period Distillation at both feature and prediction levels. (a) Multi-Scale Distillation involves downsampling the original time series into multiple coarser scales and aligning these scales between the student and teacher. (b) Multi-Period Distillation applies FFT to transform the time series into a spectrogram, followed by matching the period distributions after applying softmax.
4 Methodology
@@
$\min\nolimits_{\theta_{s}}\mathcal{L}_{sup}(\mathbf{Y},\mathbf{\hat{Y}}_{s})+\mathcal{L}_{\mathrm{KD}}^{\mathbf{Y}}(\mathbf{\hat{Y}}_{t},\mathbf{\hat{Y}}_{s})+\mathcal{L}_{\mathrm{KD}}^{\mathbf{H}}(\mathbf{H}_{t},\mathbf{H}_{s}),$ (2)
where $\theta_{s}$ is the parameter of the student; $\mathcal{L}_{sup}$ is the supervised loss (e.g., MSE) between predictions and ground truth; $\mathcal{L}_{\mathrm{KD}}^{\mathbf{Y}}$ and $\mathcal{L}_{\mathrm{KD}}^{\mathbf{H}}$ are the distillation loss terms that encourage the student model to learn knowledge from the teacher on both prediction level (Hinton, 2015) and feature level (Romero et al., 2014), respectively. Unlike conventional approaches that emphasize matching model predictions, TimeDistill integrates key time-series patterns including multi-scale and multi-period knowledge. The overall framework of TimeDistill is shown in Figure 6.
4.1 Multi-Scale Distillation

### v2

@@
Abstract.
[…]
Time Series Forecasting, Knowledge Distillation
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: the 32th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Neural networks
1. Introduction
Forecasting is a notably critical problem in the time series analysis community, which aims to predict future time series based on historical time series records (wang2024deep). It has broad practical applications such as climate modeling (wu2023interpretable), traffic flow management (yin2021deep), healthcare monitoring (kaushik2020ai) and finance analytics (granger2014forecasting).
@@ 1. Introduction
Figure 1. Performance comparison.
rise of Transformers in various domains (DBLP:journals/corr/abs-1810-04805; khan2022transformers) has led to their wide adoption in time series forecasting (wen2022transformers; autoformer; fedformer; informer; patchtst; itransformer), leveraging the strong capabilities of capturing pairwise dependencies and extracting multi-level representations within sequential data. Similarly, CNN architectures have also proven effective by developing convolution blocks for time series (moderntcn; micn). However, despite the strong performance of Transformer-based and CNN-based models, they face significant challenges in large-scale industrial applications due to their relatively high computational demands, especially in latency-sensitive scenarios like financial prediction and healthcare monitoring (granger2014forecasting; kaushik2020ai). In contrast, simpler linear or MLP models offer greater efficiency, although with lower performance (dlinear; sparsetsf). These contrasting observations raises an intriguing question:
@@
A promising approach to addressing this question is knowledge distillation (KD(hinton2015distilling), a technique that transfers knowledge from a larger and more complex model (teacher) to a smaller and simpler one (student) while maintaining comparable performance. In this work, we pioneer cross-architecture KD in time series forecasting, with MLP as the student and other advanced architectures (e.g., Transformers and CNNs) as the teacher. However, designing such a framework is non-trivial, as it remains unclear what “knowledge” should be distilled into MLP.
@@
To investigate this potential, we conduct a comparative analysis of prediction patterns between MLP and other time series models. Our findings reveal that MLP still excels on some data subsets despite its overall lower performance (Sec. 3.1), which highlights the value of harnessing the complementary capabilities across different architectures. To further explore the specific properties to distill, we focus on two key time series patterns: multi-scale pattern in temporal domain and multi-period pattern in frequency domain, given that they are vital in capturing the complex structures typical of many time series. (1) Multi-Scale Pattern: Real-world time series often show variations at multiple temporal scales. For example, hourly recorded traffic flow data capture changes within each day, while daily sampled data lose fine-grained details but reveal patterns related to holidays (timemixer). We observe that models that perform well on the finest scale also perform accurately on coarser scales, while MLP fails on most scales (Sec. 3.2). (2) Multi-Period Pattern: Time series often exhibit multiple periodicities. For instance, weather measurements may have both daily and yearly cycles, while electricity consumption data may show weekly and quarterly cycles (timesnet). We find that models that perform well can capture periodicities similar to those in the ground truth, but MLP fails to capture these periodicities (Sec. 3.2). Therefore, enhancing MLP requires distilling and integrating these multi-scale and multi-period patterns from teacher models.
@@
Based on our observations, we propose a cross-architecture KD framework named TimeDistill to bridge the performance and efficiency gap between complex teacher models and a simple MLP. Instead of solely matching predictions in conventional KD, TimeDistill focuses on aligning multi-scales and multi-period patterns between MLP and the teacher: we downsample the time series for temporal multi-scale alignment and apply Fast Fourier Transform (FFT) to align period distributions in the frequency domain. The KD process can be conducted offline, shifting heavy computations from latency-critical inference phase, where millisecond matter, to less time-sensitive training phase, where longer processing time is acceptable. We validate effectiveness of TimeDistill both theoretically and empirically and summarize our contributions as follows:
@@
Figure 2. Model efficiency comparison averaged across all prediction lengths (96, 192, 336, 720) for the ECL dataset. Full results on more datasets are listed in Appendix J.
@@
    We provide theoretical insights into benefits of TimeDistill, illustrating that proposed distillation process can be viewed as a form of data augmentation through a special mixup strategy.
@@
    We show that TimeDistill is both effective and efficient, consistently outperforming standalone MLP by up to 18.6% and surpassing teacher models in nearly all cases (see Figure 1). Additionally, it achieves up to 7x faster inference and requires up to 130× fewer parameters compared to teachers (see Figure 2).
@@
Figure 3. Win ratio (%) of MLP v.s. teacher models across datasets under input-720-predict-96 setting. The win ratio is generally large (average: 49.92%, median: 49.96%), indicating MLP and teacher models excel on different samples with minimal overlap.
@@
Figure 4. Visualization of model predictions on different downsampled scales of ECL dataset. MLP consistently shows poor performance at multiple scales, while other models perform well, highlighting the importance of capturing multi-scale patterns.
@@
Figure 5. Prediction spectrograms of various models on ECL dataset against the ground truth. MLP fails to match the amplitudes of several main frequencies in the ground truth, with red numbers indicating amplitude differences for the most significant frequency.
2. Related Work
2.1. Debate in Long-Term Time Series Forecast
The trade-off between performance and efficiency has prompted a long-standing debate between Transformer-based models and MLP in long-term time series forecasting (dlinear; lightts; sparsetsf). Informer (informer), Autoformer (autoformer), and FEDformer (fedformer) were among the leading Transformer-based methods. However, recent findings show that a simple Linear or MLP model can achieve performance comparable to these complex Transformer models across various benchmarks while offering significantly better efficiency (dlinear). This outcome has raised questions about the necessity of Transformers in time series forecasting. Following this, research has moved in two directions. One direction suggests that the issues with Transformer-based models arise from the way they are applied. For example, PatchTST (patchtst) uses patching to preserve local information, and iTransformer (itransformer) focuses on capturing multivariate correlations. These approaches surpass the simple one-layer MLP and demonstrate that Transformers can still deliver strong results in time series forecasting if applied effectively. Meanwhile, CNN-based models have also shown strong performance similar to Transformer-based models. TimesNet (timesnet) transforms 1D time series into 2D variations and applies 2D CNN kernels, MICN (micn) adopt multi-scale convolution structures to capture local features and global correlations, and ModernTCN (moderntcn) proposes a framework with much larger receptive fields than prior CNN-based structures. Nevertheless, these powerful CNN-based models also face efficiency issues, which further broadens the scope of the debate between performance and efficiency. The other direction focuses on developing lightweight MLP, such as N-BEATS (nbeats), N-hits (nhits), LightTS (lightts), FreTS (frets), TSMixer (tsmixer), TiDE (tide), which offer improved efficiency. However, these models typically only match, rather than surpass, the performance of state-of-the-art Transformer-based methods and CNN-based methods. In summary, while Transformer-based and CNN-based models generally offer better performance, simple MLP is more efficient. Therefore, we managed to combine the performance of Transformer-based and CNN-based models with MLP to produce a powerful and efficient model by cross-architecture KD.
2.2. Knolwedge Distillation on Time Series
Knowledge distillation (KD(hinton2015distilling) transfers knowledge from a larger, more complex model (teacher) to a smaller, simpler model (student) while maintaining comparable performance. By aligning the output distributions of teacher and student models, KD provides richer training signals than hard labels alone, enabling the student to capture subtle patterns that the teacher has learned. In the context of time series, CAKD (xu2022contrastive) introduces a two-stage distillation scheme that distills features using adversarial and contrastive learning and performs prediction-level distillation. LightTS (campos2023lightts) designs a KD framework specifically for cases where the teacher is an ensemble classifier in time series classification tasks, which limits its applicability to teachers with other architectures. Both of these works do not incorporate time series-specific designs. In contrast, our framework emphasizes extracting essential time series patterns, including multi-scale and multi-period patterns, enabling more effective knowledge distillation. Additionally, we are the first to explore cross-architecture KD.
3. Preliminary Studies
@@
3.1. What Motivates Distillation?
@@
While MLP excel in efficiency, they often lag in performance compared to Transformer and CNN models or achieve only similar results (Figure 1). However, even if MLP underperforms its teachers overall, it may excel on specific samples. To investigate this, we compare prediction errors of MLP’s prediction $\mathbf{\hat{Y}}_{s}$ and teacher model’s prediction $\mathbf{\hat{Y}}_{t}$ for $N$ samples, defined as $e_{s}=\{\text{MSE}(\mathbf{\hat{Y}}_{s}^{1},\mathbf{Y}^{1}),\cdots,\\ \text{MSE}(\mathbf{\hat{Y}}_{s}^{N},\mathbf{Y}^{N})\}$ and $e_{t}=\{\text{MSE}(\mathbf{\hat{Y}}_{t}^{1},\mathbf{Y}^{1}),\cdots,\text{MSE}(\mathbf{\hat{Y}}_{t}^{N},\mathbf{Y}^{N})\}.$ We calculate win ratio, indicating where MLP outperforms teachers:
(1) $\text{Win Ratio}={\sum\mathbbm{1}(e_{s}<e_{t})}/{N},$
@@
where $\mathbbm{1}(\cdot)$ equals 1 if MLP outperforms the teacher, otherwise 0. As shown in Figure 5, the win ratio is high (average: 49.92%, median: 49.96%), despite MLP underperforming teacher models overall. For example, on the Traffic dataset, MLP lags behind ModernTCN but wins on 81.19% of the samples, indicating differing strengths on distinct subsets. This highlights the potential of distilling complementary knowledge from the teacher model into MLP.
@@
An intuitive strategy of KD for MLP is to align predictions with teachers, but this faces limitations. First, it risks overfitting noise in the teacher’s predictions, leading to less stable knowledge (chen2017learning; takamoto2020efficient; gajbhiye2021knowledge). Second, MLP may struggle to replicate the complex patterns such as seasonality, trends, and periodicity in teacher predictions directly due to their limited capacity. Third, this approach overlooks valuable knowledge from intermediate features of teacher model. Therefore, the specific knowledge to distill into MLP requires further exploration.
3.2. What Should We Distill?
@@
To investigate which complementary time series patterns to distill into MLP, we analyze prediction patterns in both temporal and frequency domains, considering real-world variations across temporal scales (timemixer) and periodicities (timesnet). Thus, we conduct further analysis of these two patterns by presenting illustrative cases. The implementation is detailed in Appendix E. (1) Multi-Scale Pattern: By downsampling predictions $\mathbf{\hat{Y}}$ using convolutional operations, we obtain their multi-scale representations. Figure 5 shows the analysis of multi-scale temporal patterns by downsampling the time series (Scale 0) to coarser scales (Scales 1–3). Models excelling at Scale 0 generally perform well across all scales. The teacher models align closely with the ground truth at Scale 3, capturing underlying trends effectively. In contrast, MLP significantly deviates, showing its limitations in handling multi-scale patterns. (2) Multi-Period Pattern: Periodicities in time series are visible in the frequency domain by transforming time series into spectrograms, with the x-axis representing frequency and periodicity calculated as the time series length $S$ divided by frequency. Figure 5 shows spectrograms of model predictions and the ground truth. Models with lower MSE capture periodicities more accurately, closely matching the ground truth at dominant frequencies. In contrast, MLP predictions show larger discrepancies, highlighting its inability to effectively capture periodic patterns.
@@
From our observations, we conjecture that MLP underperforms on certain samples due to their inability to capture essential multi-scale and multi-period patterns, which are essential for effective time series forecasting. To improve MLP, it is crucial to incorporate these complementary patterns from teachers during distillation.
@@
Figure 6. Overall framework of TimeDistill, which distills knowledge from a teacher model to a student MLP using (a) Multi-Scale Distillation and (b) Multi-Period Distillation at both feature and prediction levels. (a) Multi-Scale Distillation involves downsampling the original time series into multiple coarser scales and aligning these scales between the student and teacher. (b) Multi-Period Distillation applies FFT to transform the time series into a spectrogram, followed by matching the period distributions after applying softmax.
4. Methodology
@@
(2) $\min\nolimits_{\theta_{s}}\mathcal{L}_{sup}(\mathbf{Y},\mathbf{\hat{Y}}_{s})+\mathcal{L}_{\mathrm{KD}}^{\mathbf{Y}}(\mathbf{\hat{Y}}_{t},\mathbf{\hat{Y}}_{s})+\mathcal{L}_{\mathrm{KD}}^{\mathbf{H}}(\mathbf{H}_{t},\mathbf{H}_{s}),$
where $\theta_{s}$ is the parameter of the student; $\mathcal{L}_{sup}$ is the supervised loss (e.g., MSE) between predictions and ground truth; $\mathcal{L}_{\mathrm{KD}}^{\mathbf{Y}}$ and $\mathcal{L}_{\mathrm{KD}}^{\mathbf{H}}$ are the distillation loss terms that encourage student model to learn knowledge from teacher on both prediction level (hinton2015distilling) and feature level (romero2014fitnets). Unlike conventional approaches that emphasize matching model predictions, TimeDistill integrates key time-series patterns including multi-scale and multi-period knowledge. The overall framework of TimeDistill is shown in Figure 6.
4.1. Multi-Scale Distillation